For sound waves in air with frequency 1000 Hz a displacement amplitude of 1.2 \times 10^{ - 8} \;{\rm{ m}} produces a pressure amplitude of 3.0 \times 10^{ - 2}\; {\rm{ Pa}}. Use v_{\rm sound}= 344 {\rm m/s}.

a) What is the wavelength of these waves?

b) For 1000-Hz waves in air, what displacement amplitude would be needed for the pressure amplitude to be at the pain threshold, which is 30 Pa?

Express your answer using two significant figures.
c) For what wavelength will waves with a displacement amplitude of 1.2 \times 10^{ - 8}\; {\rm{ m}} produce a pressure amplitude of 1.5 \times 10^{ - 3}\; {\rm{ Pa}}?
Express your answer using two significant figures.
d) For what frequency will waves with a displacement amplitude of 1.2 \times 10^{ - 8}\; {\rm{ m}} produce a pressure amplitude of 1.5 \times 10^{ - 3}\; {\rm{ Pa}}?
Express your answer using two significant figures.
Answer #1
Concepts and reason

The concepts required to solve this problem is wave number, wave speed, and pressure amplitude equation of the sound wave.

Initially, calculate the wavelength of the wave by using the expression of the wave speed. Calculate the wavenumber of the wave and then calculate the value of the bulk modulus by using the pressure amplitude equation of the wave. After that, calculate the displacement amplitude of the wave by using the pressure amplitude equation of the wave.

Later, calculate the wavelength of the wave by using the expression for wave number.

Finally, calculate the frequency of the wave by using wave speed.

Fundamentals

The expression for wave speed is,

v=λfv = \lambda f

Here, f is the frequency and λ\lambda is the wavelength.

The expression for wave number is,

k=ωvk = \frac{\omega }{v}

Here, ω\omega is the angular frequency and v is the speed of the sound wave.

The expression for angular frequency is,

ω=2πf\omega = 2\pi f

Here, f is the frequency.

The expression for the pressure amplitude of the wave is,

P=BkAP = BkA

Here, B is the bulk modulus, k is the wave number, and A is the displacement amplitude of the wave.

(a)

The expression for wave speed is,

v=λfv = \lambda f

Rewrite the above expression for λ\lambda .

λ=vf\lambda = \frac{v}{f}

Substitute 344 m/s for v and 1,000 Hz for f in the above equation.

λ=344m/s1,000Hz=0.344Hz\begin{array}{c}\\\lambda = \frac{{344\;{\rm{m/s}}}}{{1,000\;{\rm{Hz}}}}\\\\ = 0.344\;{\rm{Hz}}\\\end{array}

Therefore, the wavelength of these waves is 0.344 Hz.

(b)

The expression for wave number is,

k=ωvk = \frac{\omega }{v}

The expression for angular frequency is,

ω=2πf\omega = 2\pi f

Substitute 2πf2\pi f for ω\omega in the equation k=ωvk = \frac{\omega }{v}.

k=2πfvk = \frac{{2\pi f}}{v}

Substitute 344 m/s for v and 1,000 Hz for f in the above equation.

k=2(3.14)(1,000Hz)(344m/s)=18.25rad/m\begin{array}{r}\\k = \frac{{2\left( {3.14} \right)\left( {1,000\;{\rm{Hz}}} \right)}}{{\left( {344\;{\rm{m/s}}} \right)}}\\\\ = 18.25\;{\rm{rad/m}}\\\end{array}

The expression for the pressure amplitude of the wave is,

P=BkAP = BkA

Rewrite the above expression for B.

B=PkAB = \frac{P}{{kA}}

Substitute 3.0×102Pa3.0 \times {10^{ - 2}}\;{\rm{Pa}} for P , 18.25 rad/m for k, and 1.2×108m1.2 \times {10^{ - 8}}\;{\rm{m}} for A in the above equation.

B=3.0×102Pa(18.25rad/m)(1.2×108m)=1.369×105Pa\begin{array}{c}\\B = \frac{{3.0 \times {{10}^{ - 2}}\;{\rm{Pa}}}}{{\left( {18.25\;{\rm{rad/m}}} \right)\left( {\;1.2 \times {{10}^{ - 8}}\;{\rm{m}}} \right)}}\\\\ = 1.369 \times {10^5}\;{\rm{Pa}}\;\\\end{array}

Calculate the displacement amplitude of the wave as follows:

P=BkAP = BkA

Rewrite the above expression for A.

A=PBkA = \frac{P}{{Bk}}

Substitute 3.0×102Pa3.0 \times {10^{ - 2}}\;{\rm{Pa}}30 Pa for P , 18.25 rad/m for k, and 1.369×105Pa1.369 \times {10^5}\;{\rm{Pa}} for B in the above equation.

A=3.0×102Pa(1.369×105Pa)(18.25rad/m)=1.2×105m\begin{array}{c}\\A = \frac{{3.0 \times {{10}^{ - 2}}\;{\rm{Pa}}}}{{\left( {1.369 \times {{10}^5}\;{\rm{Pa}}} \right)\left( {18.25\;{\rm{rad/m}}} \right)}}\\\\ = 1.2 \times {10^{ - 5}}\;{\rm{m}}\\\end{array}

Therefore, the displacement amplitude of the wave is 1.2×105m1.2 \times {10^{ - 5}}\;{\rm{m}}.

(c)

Calculate the wave number as follows:

P=BkAP = BkA

Rewrite the above expression for k.

k=PBAk = \frac{P}{{BA}}

Substitute 1.5×103Pa1.5 \times {10^{ - 3}}\;{\rm{Pa}} for P, 1.369×105Pa1.369 \times {10^5}\;{\rm{Pa}} for B, and 1.2×108m1.2 \times {10^{ - 8}}\;{\rm{m}} for A in the above equation.

k=(1.5×103Pa)(1.369×105Pa)(1.2×108m)=0.91rad/m\begin{array}{c}\\k = \frac{{\left( {1.5 \times {{10}^{ - 3}}\;{\rm{Pa}}} \right)}}{{\left( {1.369 \times {{10}^5}\;{\rm{Pa}}} \right)\left( {1.2 \times {{10}^{ - 8}}\;{\rm{m}}} \right)}}\\\\ = 0.91\;{\rm{rad/m}}\\\end{array}

The expression for wavelength is,

λ=2πk\lambda = \frac{{2\pi }}{k}

Substitute 0.91 rad/m in the above equation.

λ=2(3.14)(0.91rad/m)=6.9m\begin{array}{c}\\\lambda = \frac{{2\left( {3.14} \right)}}{{\left( {0.91\;{\rm{rad/m}}} \right)}}\\\\ = 6.9\;{\rm{m}}\\\end{array}

Therefore, the wavelength is 6.9 m.

(d)

The expression for frequency is,

f=vλf = \frac{v}{\lambda }

Substitute 344 m/s for v and 6.9 m for λ\lambda in the above equation.

f=344m/s6.9m=49.8550Hz\begin{array}{c}\\f = \frac{{344\;{\rm{m/s}}}}{{6.9\;{\rm{m}}}}\\\\ = 49.85\\\\ \approx 50\;{\rm{Hz}}\\\end{array}

Therefore, the required frequency is 50 Hz.

Ans: Part a

Therefore, the wavelength of these waves is 0.344 Hz.

Part b

Therefore, the displacement amplitude of the wave is 1.2×105m1.2 \times {10^{ - 5}}\;{\rm{m}}.

Part c

Therefore, the wavelength is 6.9 m.

Part d

Therefore, the required frequency is 50 Hz.